OpenStudy (anonymous):

f(x)=1/x-2+1 Find x if f(x) = -1

OpenStudy (anonymous):



hero (hero):

Hint: Since f(x) = -1, then: \(\frac{1}{x-2} + 1 = -1\)

OpenStudy (anonymous):

so it's -1

hero (hero):

You have to solve for x obviously.

OpenStudy (anonymous):

so x=1

OpenStudy (anonymous):

so what is f(x) of -1

OpenStudy (anonymous):

-1=1/x-2+1 x+2-1=1+1 x+2=2+1 x=2+1-2 =1

OpenStudy (anonymous):

hence x is 1

OpenStudy (anonymous):

@Kingjames1 u r cleared nw

OpenStudy (anonymous):

ok f(x)=1

OpenStudy (anonymous):

we wanted to x know

OpenStudy (anonymous):

then we want to write as x=1